# Partial confinement in a quantum-link simulator

### v1

@@ VI Conclusion
To conclude, we have presented a comprehensive investigation of partial confinement on the platform of the spin-1 quantum link model, a promising platform realizable with cold atoms in optical superlattices. The partial confinement is characterized by a dependency of confinement properties on the spatial arrangement of charged particles, manifested by the asymmetry of the equilibrium energy and the string tension of the string state with respect to the charge ordering. In the non-equilibrium dynamics, string states with different charge orderings exhibit strikingly distinct dynamical features for such observables as local fermion occupations, bipartite entanglement entropy, and edge charge correlations. We have also elucidated that manipulating the topological angle can be an effective proxy for controlling charge ordering, thereby simplifying experimental procedures by obviating the need for direct charge manipulation. Given that the quantum link model is amenable to current experimental capabilities, our study offers a strategic avenue for exploring novel physics in gauge theories using state-of-the-art quantum simulators.

### v2

@@ VI Conclusion
To conclude, we have presented a comprehensive investigation of partial confinement on the platform of the spin-1 quantum link model, a promising platform realizable with cold atoms in optical superlattices. The partial confinement is characterized by a dependency of confinement properties on the spatial arrangement of charged particles, manifested by the asymmetry of the equilibrium energy and the string tension of the string state with respect to the charge ordering. In the non-equilibrium dynamics, both string and meson states exhibit strikingly distinct dynamical features depending on their (de)confining status, as reflected in such quantities as local fermion occupations, bipartite entanglement entropy, and edge charge correlations. We have also elucidated that manipulating the topological angle can be an effective proxy for controlling charge ordering, thereby simplifying experimental procedures by obviating the need for direct charge manipulation. Given that the quantum link model is amenable to current experimental capabilities, our study offers a strategic avenue for exploring novel physics in gauge theories using state-of-the-art quantum simulators. Furthermore, it may also be interesting to discuss the partial confinement in other forms of lattice gauge theories such as the improved Hamiltonian [91, 92].